# Electron Inertia and Magnetic Reconnection

### v4

@@ II.2 Simplified form for $\vec{\mathcal{E}}$
The derivation of the simplified form starts with replacing $\vec{B}$ by $\vec{\mathcal{B}}$ the Hall term.
@@ IV Chaos and the applicability of the non-reconnection of $\vec{\mathcal{B}}$
Magnetic field line chaos fundamentally changes the theory of magnetic reconnection. The effect of chaos on election-inertia-produced reconnection in specific magnetic geometries was investigated in four papers by Borgogno and collaborators Borgogno:2005 ; Borgogno:2008 ; Borgogno:2011a ; Borgogno:2011b from 2005 to 2011.
@@
A general mathematical theory of chaos and magnetic reconnection involves concepts such a cantori and turnstiles, which are not broadly known in the plasma physics community but were recently reviewed in Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas, Boozer:chaos . Simple derivations of effects associated with magnetic reconnection were given in the 2025 article, Magnetic reconnection and dynamos in the presence of plasma turbulence, Boozer:2025 . In particular, it is shown that a chaotic flow of the magnetic field lines causes an exponential increase in the magnetic field energy in two-dimensional systems—the magnetic field lines themselves cannot become chaotic. In three-coordinate systems, a chaotic field-line flow gives chaotic magnetic field lines.
@@ Acknowledgements
The author would like to thank Peter Constantin, the John von Neumann Professor of Mathematics at Princeton University, who was sufficiently intrigued by the result that he sent me an alternative derivation. The author would also like to thank Amitava Bhattacharjee, who suggested reference Ham-recon:1998 , and Andrew Brown, who pointed out reference Inertial-MHD:2015 . Most importantly, he would like to thank Luca Comisso, who pointed out the relativistic derivation of $\vec{\mathcal{B}}$ though he did so relatively late, December 24, 2025.

### v5

@@ II.2 Simplified form for $\vec{\mathcal{E}}$
The derivation of the simplified form starts with replacing $\vec{B}$ by $\vec{\mathcal{B}}$ in the Hall term.
@@ II.3 Reconnection condition
When the magnetic field lines are carried by a flow $\vec{u}_{\bot}$, field lines that lie in separate tubes of magnetic flux remain separated forever. But as illustrated in Figure 1, the surfaces bounding tubes of magnetic flux can become extremely contorted, which makes the preservation of distinct flux tubes hypersensitive to diffusive effects such resistivity. See the discussion of Equation (37).
@@ IV Chaos and the applicability of the non-reconnection of $\vec{\mathcal{B}}$
Magnetic field line chaos fundamentally changes the theory of magnetic reconnection, Figure 1. The effect of chaos on election-inertia-produced reconnection in specific magnetic geometries was investigated in four papers by Borgogno and collaborators Borgogno:2005 ; Borgogno:2008 ; Borgogno:2011a ; Borgogno:2011b from 2005 to 2011. In their 2005 paper, Borgorno et al carried out a numerical study of collisonless but unstable plasmas that initially had nested magnetic surfaces. This means spatially bounded constant-$\psi(\vec{x},t)$ surfaces with $\vec{B}\cdot\vec{\nabla}\psi=0$. If the magnetic evolution were perfectly ideal, the constant-$\psi(\vec{x},t)$ surfaces could be deformed but not destroyed. Nevertheless, the weak non-ideal effects present in a collisionless plasma can cause these surfaces to break and create spatial volumes in which the magnetic field lines become chaotic. In volumes of space that contain both constant-$\psi(\vec{x},t)$ surfaces and chaotic regions, Borgorno et al Borgogno:2005 measured the reconnected flux by the sum of the magnetic fluxes in all of the chaotic regions that are bounded by the remaining constant-$\psi(\vec{x},t)$ surfaces. This reconnected flux was found to increase exponentially in time at approximately the instability growth rate once the instability amplitude was sufficient to produce chaotic regions. Borgogno et al stated in their paper Borgogno:2011a : “The development of magnetic field line reconnection in three-dimensional configurations is an intrinsically chaotic process.” Although true, this is recognized by only a remarkably small fraction of those studying magnetic reconnection.
@@
A general mathematical theory of chaos and magnetic reconnection involves concepts such a cantori and turnstiles, which are not broadly known in the plasma physics community but were recently reviewed in Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas, Boozer:chaos . Cantori, although not by that name, were discussed in two papers Borgogno:2011a ; Borgogno:2011b published by Borgorno et al in 2011. They noted that “certain surfaces may act as barriers to this magnetic field line penetration,” the penetration of magnetic field lines through surfaces near which they linger.
Simple derivations of effects associated with magnetic reconnection were given in the 2025 article, Magnetic reconnection and dynamos in the presence of plasma turbulence, Boozer:2025 . In particular, it is shown that a chaotic flow of the magnetic field lines causes an exponential increase in the magnetic field energy in two-dimensional systems—the magnetic field lines themselves cannot become chaotic. In three-coordinate systems, a chaotic field-line flow gives chaotic magnetic field lines.
@@ Acknowledgements
The author would like to thank Peter Constantin, the John von Neumann Professor of Mathematics at Princeton University, who was sufficiently intrigued by the result that he sent me an alternative derivation. The author would also like to thank Amitava Bhattacharjee, who suggested reference Ham-recon:1998 , and Andrew Brown, who pointed out reference Inertial-MHD:2015 . Most importantly, he would like to thank Luca Comisso, who pointed out the relativistic derivation of $\vec{\mathcal{B}}$.